with the current, joe can row a canoe 3 miles in 12 mins. Against the current, he requires 18 minutes to row same distance. HOw fast does joe row in still water, and how fast is the current?
Answer by nyc_function(2733)

Let x = Joe's speed in still water
Let y = speed of current
Since there are two things we are solving for, we are looking at a system of linear equations in two variables.

Against Current:
rate = x - y
time = 0.20
distance = 3

0.20(x - y) = 3
0.20x - 0.20y = 3...This is our first equation.
===============================================

With Current:
rate = x + y
time = 0.30
distance = 3
0.30(x + y) = 3
0.30x + 0.30y = 3...This is our second equation.
================================================
Here they are together:
0.20x - 0.20y = 3
0.30x + 0.30y = 3
Can you finish now?